OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = −4 5x + 3y = −6

Nnesha (nnesha):

you can do elimination method substitution graphing matrices so which one is easy for u??? please let me know :)

OpenStudy (anonymous):

um

Nnesha (nnesha):

so i guess i should select is that okay ??

OpenStudy (anonymous):

oh I just got it using a graph, sorry for wasting ur time.

Nnesha (nnesha):

ohh okay so what is y ??

OpenStudy (anonymous):

-1,-2

OpenStudy (dtan5457):

you probably wanna use substitution. it's the easiest for here. 2x+y=-4 5x+3y=-6 isolate y in the first equation. y=-4-2x plug that into the 2nd equation 5x+3(-4-2x)=-6 that is 5x-12-6x=-6 that equates to -x=6, which means x=-6. to get y, plug back into first equation. 2(-6)+y=-4 -12+y=-4 add 12, y=8

OpenStudy (anonymous):

okay, that makes a lot of sense more then the graph my brother tried to show me....

OpenStudy (dtan5457):

graphing is actually really easy if you know how to do it, but if you don't have that in handy...do it algebraically.

OpenStudy (anonymous):

so x = -6 and y = 8 ?

OpenStudy (dtan5457):

yes.

OpenStudy (anonymous):

thanks, you get a medal.

OpenStudy (anonymous):

can you help with another one?

OpenStudy (dtan5457):

I'm sure @Nnesha will like to give it a try

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The graph of f(x) = (0.5)x is replaced by the graph of g(x) = (0.5)x - k. If g(x) is obtained by shifting f(x) down by 5 units, the value of k is

Nnesha (nnesha):

graphing method i don't how you did but its really easy point where both lines intersect i graphed it 4 u

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

\(\huge\color{blue}{my ~ pleasure }\)